find the center of circle with these three points ( -1,3) (-13,3) (-7,9)
Answer by Alan3354(69443) 
 find the center of circle with these three points (-1,3) (-13,3) (-7,9)
-----------------
A(-1,3) B(-13,3) C(-7,9)
---
X Y x^2+y^2 Coeff
Point 1 -1 3 10 1
Point 2 -13 3 178 1
Point 3 -7 9 130 1
=========================
Using the matrix above, the center is at (-7,3)
-72(x^2+y^2) - 1008x + 432y = 1584
x^2 + 14x + y^2 - 6y + 72 = 0
(x+7)^2 + (y-3)^2 = 36
|
|
|